OpenStudy (anonymous):

Choose the correct simplification of the expression c2 • c9. answer is c18

OpenStudy (anonymous):

I think...

OpenStudy (anonymous):

no Don't multiply 2 and 9 Add them It should be c^11

OpenStudy (solomonzelman):

c^2 × c^9 = c^(2+9 ) = c^11

OpenStudy (anonymous):

ooo,i get it now,thanks Bobmath

OpenStudy (anonymous):

yw